### v1

@@ 2.3 Low frequency noise reconstruction
1/f noise represents the spectral-domain manifestation of multiple RTN defects fluctuating simultaneously, with time constants following a log-normal distribution. Indeed, the power spectral density (PSD) of a single defect RTN has a Lorenztian shape, and summing the contribution of many defects whose time constants are log-normally distributed results in a 1/f spectrum [40]. Based on this consideration, PSD of converted single raw traces were calculated and summed over. Afterwards, the sum of all the PSDs has been divided by the total number of device measured. The LFN spectra obtained must be regarded as the noise of the entire device ensemble. In variability studies, the geometric (or logarithmic) mean is sometimes used to estimate the average noise, as it is less sensitive to the performance of devices in the tails of the statistical distribution, i.e., outliers [40, 41]. In the present work, however, the aim is to investigate the contribution to the 1/f noise of all detected defects, including those located in the tail of the trimodal exponential distribution shown in Fig. 2(a).
@@
Before discussing the details of the LFN analysis, it is worth to point out that the noise coming from both the digital circuitry of the array and the experimental setup does not affect the reconstructed 1/f spectra substantially. To verify that, the noise calculated by taking into account just those traces which showed RTN was compared with the noise obtained using all the measured traces, which are almost three times more. In Fig. 4(a) is shown the noise power for sake of clarity. The noise level does not change significantly, meaning that the 1/f spectra are dominated by the detected devices’ RTN. Furthermore, Fig. 4(a) also provides a positive feedback on the RTN extraction method previously applied, showing that most of defects activity has been correctly detected.
@@
The input-referred 1/f drain current noise s follow a 1/f behaviour (see Fig. 4(b)), consistent with the commonly accepted theories, and they increase at cryogenic temperatures, as widely reported in literature [10, 11, 13, 12]. The value of extracted at Hz is plotted in Fig. 4(c) as function of the overdrive. At V, is higher at low temperatures because of both larger percolation and the activity of defects belonging to the large branch, as discussed later. Increasing the gate voltage, the channel gets stronger, and at cryogenic temperatures decreases, which is consistent with the vanishing of the third branch in the CCDF (see Fig. 3(a)). Above V, slightly increases, especially at K. This trend could be attributed to a shift in the defect active energy area scanned by the channel Fermi level. Indeed, as increases, the Fermi level probes an increasing number of defects in the layer, while traps in the move outside the energy window swept by the Fermi level. This interpretation is supported by the modality of the CCDFs discussed in the previous section, which become monomodal above V, and also by the trend in Fig. 4(d), where the contribution to of defects associated to different modes is investigated. At V and K, bulk defects still play a significant role in the LFN. Indeed, an examination of Fig. 4(c) and Fig. 4(d) reveals that the component related to layer is comparable to, or even higher than, observed at K. This finding is consistent with the analysis discussed in the previous section, where it was shown that more than 80% of the total RTN activity came from the layer. Nevertheless, it is worth to investigate further the implications of this result.
@@
Since the gate voltages considered here are around , LFN is assumed to be in the carrier number fluctuation (CNF) regime. can then be written as
@@
where is the Boltzmann constant, is the volumetric gate dielectric trap density per unit energy, is the oxide capacitance per unit area, is the frequency and is the wentzel-kramers-brillouin (WKB) tunnelling probability for a rectangular barrier. The factor in eq.(3) originates from the integral , being the trap energy and the near-equilibrium occupation function of the defect, which is described by a Fermi function with Fermi level [10]. The integral above is then the area of the bell-shaped function centred at (see Fig. 5(a)). In other words, only defects with energies within a few of the Fermi level can contribute to the noise. This implies that, as shown in Fig. 5(a), at K the active energy area is almost two orders of magnitude smaller than at K. On the other hand, the total number of defects, extracted by fitting distributions, which contributes to RTN and, therefore, to LFN, increases dramatically at cryogenic temperatures (see Fig. 5(b)). One possible way to reconcile these results is to postulate that the defect density increases at low temperatures. As already stressed, at V, most of active defects are in the layer. Hence, the following discussion focuses on the physical origin of electron trapping in this material. Several works [42, 43, 44, 45] showed that electron trapping in is related to disorder, which creates electron traps associated with low-coordinated sites and elongated bonds. The trapping is followed by significant polaron-like distortion of the surrounding amorphous network. At cryogenic temperatures, due to the limited thermal energy available, an electron has a significantly higher probability of remaining near the precursor site in the near-interface oxide long enough to polarize the oxide network, resulting in a shallow trapping event. Afterwards, the electron can either tunnel into this self-created trap state or remain nearby without contributing to the channel current.
@@
defects are the main responsible for the high RTN activity at cryogenic temperatures, but their contribution to LFN is the smallest if compared to the defects of the other branches (see Fig. 4(d)). Indeed, traps associated to the second ( layer) and third mode contribute almost equally to . As a result, the layer hosts the majority of the defects active at cryogenic temperatures. However, the dominant contribution to LFN still arises from interface traps which -according to the previous discussion- include defects associated with both the second and third branches of the CCDF in Fig. 2(a). This opposite trend arises from the larger of the interface traps, which is proportional to the amplitude of the PSD of the original RTN trace. As already discussed, at V, is primarily associated with defects due to shift in the active energy area swept by the Fermi level. The physical origin of the kinetics of these defects, which do not freeze out, is still under debate. Recently, in was shown that a full quantum-mechanical solution of the non-radiative multi-phonon (NMP) model can interpret the experimental results at K [18, 9, 46]. According to this solution, nuclear tunnelling becomes the predominant mechanism of charge trapping at cryogenic temperatures. Beside that, the number of detected defects in the layer at higher overdrives is weakly temperature dependent (see Supplement Material), which is challenging to reconcile with the previous discussion based on the trend shown if Fig. 5(a). Further analyses are required to comprehend the nature of charge trapping in at cryogenic temperatures.
@@ 2.4 Time constants and step heights correlation
Since determines the defect amplitude, if charge trapping processes are dominated by elastic tunnelling, there must be a correlation between and . Many works already proved the inadequacy of elastic tunnelling and, therefore, of McWhorter model to describe LFN in MOSFET devices at K [19, 48, 49]. Here, the possible correlation between and is investigated from K down to K. Details about the extraction of time constants are provided in Supplement Material. It is worth to stress that, since the depends with the trap position and, according to eq.(4), depends with the trap position, a linear correlation on a semi-logarithmic axis is expected. In 6(a) and 6(b) the scatter plots at K are reported, together with the calculated Pearson coefficient . No correlation is found, in agreement with [19, 48, 49]. Decreasing temperature, results are similar (see Supplement Materials), down to K (6(c) and 6(d)). From these results it is evident that McWhorter model is not accurate in describing LFN and that defect kinetic is not dominated by the elastic tunnelling, neither at cryogenic temperatures.
@@ 3 Conclusion
The first part of the paper was devoted to the analysis of the threshold voltage shifts due to charged defects, reported in cumulative plots. In this context, it was shown that, at K, the CCDF followed a monomodal exponential distribution, and only defects associated to layer contributed to RTN. At cryogenic temperatures, instead, the CCDF changed its modality with the appearing of other two modes: a small one, attributed to defects, and a large one whose origin is still controversial. An interpretation of both the trimodal shape of the CCDF and the nature of the third branch was provided exploiting the percolation theory. The change in the shape of the CCDF at cryogenic temperatures was attributed to a major sensitivity of the inversion carriers to channel inhomogeneities. Indeed, even though random dopants and surface potential fluctuations are the main cause of the exponential distribution of the already at K, at low temperature carriers’ thermal energy is reduced and they are more susceptible to channel disorder. This larger non-uniformity can induce the formation of insulating regions in the channel, increasing the electrostatic impact of oxide and interface defects: this could origin the third mode of large experimentally observed. Nonetheless, the primary outcome of the first part of this study is the identification of electron trapping within the layer as the dominant contributor for over 80% of the total RTN measured at K.
The second half of this work is dedicated to low frequency noise. The aim is to investigate the contribution of individual defect branches to the total 1/f noise, reconstructed using the original time traces. Even though most of active defects at cryogenic temperatures are associated to the layer, their electrostatic impact is quite small if compared to interface traps. Therefore, most of low frequency noise arises from interfaces defects. Few defects cause most of the 1/f noise at cryogenic temperatures. Nevertheless, LFN generated by bulk traps at K is still comparable to the total LFN at K. An interpretation of this huge activity of bulk defects was proposed and it was based on the concept of electron self-localization due to the polarization of the network. With increasing overdrive voltage, the noise contribution from traps steadily diminishes until it becomes negligible.
@@ 4.1 Smart Array
The chip design also incorporates sense lines at the output of the shift register, as well as on the gate and drain of a device within the array. This allows us to verify the functionality of the periphery and identify potential voltage drops on the gate and drains. Moreover, all measurements are performed using a custom-built switching matrix with separately controllable voltage sources to manage the periphery, along with K2600 SMUs for current readouts.

### v2

@@ 2.3 Low frequency noise reconstruction
1/f noise represents the spectral-domain manifestation of multiple RTN defects fluctuating simultaneously, with time constants following a log-normal distribution. Indeed, the power spectral density (PSD) of a single defect RTN has a Lorenztian shape, and summing the contribution of many defects whose time constants are log-normally distributed results in a 1/f spectrum [42]. Based on this consideration, PSD of converted single raw traces were calculated and summed over. Afterwards, the sum of all the PSDs has been divided by the total number of device measured. The LFN spectra obtained must be regarded as the noise of the entire device ensemble. In variability studies, the geometric (or logarithmic) mean is sometimes used to estimate the average noise, as it is less sensitive to the performance of devices in the tails of the statistical distribution, i.e., outliers [42, 43]. In the present work, however, the aim is to investigate the contribution to the 1/f noise of all detected defects, including those located in the tail of the trimodal exponential distribution shown in Fig.˜2(a).
@@
Before discussing the details of the LFN analysis, it is worth to point out that the noise coming from both the digital circuitry of the array and the experimental setup does not affect the reconstructed 1/f spectra substantially. To verify that, the noise calculated by taking into account just those traces which showed RTN was compared with the noise obtained using all the measured traces, which are almost three times more. In Fig.˜4(a) is shown the noise power for sake of clarity. The noise level does not change significantly, meaning that the 1/f spectra are dominated by the detected devices’ RTN. Furthermore, Fig.˜4(a) also provides a positive feedback on the RTN extraction method previously applied, showing that most of defects activity has been correctly detected.
@@
The input-referred 1/f drain current noise follows a 1/f behaviour (see Fig.˜4(b)), consistent with the commonly accepted theories, and it increases at cryogenic temperatures, as widely reported in literature [10, 11, 13, 12]. The value of extracted at Hz is plotted in Fig.˜4(c) as function of the overdrive. At V, is higher at low temperatures because of both larger percolation and the activity of defects associated with the large branch, as discussed later. Increasing the gate voltage, the channel gets stronger, and at cryogenic temperatures decreases, which is consistent with the vanishing of the third branch in the CCDF (see Fig.˜3(c)). At V, slightly increases, especially at K. This trend could be attributed to a shift in the defect active energy area scanned by the channel Fermi level. Indeed, as increases, the Fermi level probes an increasing number of defects in the layer, while traps in the move outside the energy window swept by the Fermi level. This interpretation is supported by the modality of the CCDFs discussed in the previous section, which become monomodals at V, and also by the trend in Fig.˜4(d), where the contribution to of defects associated to different modes is investigated. At V and K, bulk defects still play a significant role in the LFN. Indeed, an examination of Fig.˜4(c) and Fig.˜4(d) reveals that the component related to layer is comparable to, or even higher than, the total observed at K. This finding is consistent with the analysis discussed in the previous section, where it was shown that more than 80% of the total RTN activity came from the layer. By inspection of Fig.˜2(c) and Fig.˜4(d), one can conclude that defects are the main responsible for the high RTN activity at cryogenic temperatures and at V, but their contribution to LFN is small if compared to that of defects associated with the other branches, which contribute almost equally to the total . This opposite trend is consequence of the fact that the amplitude of the PSD of the original RTN traces is directly proportional to , which is larger for the interface traps. At V, the CCDF is still trimodal, but the contribution to the LFN of the third mode is mitigated. At V and V, most of the noise is attributed to defects, whereas the contribution of the branch is almost unchanged from V. At V, CCDFs are monomodal and is fully associated with traps in the band.
@@
It is worth to further investigate some consequences of the previous discussions on the possible oxide defects origin. Since the gate voltages considered here are around , LFN is assumed to be in the carrier number fluctuation (CNF) regime. can then be written as
@@
where is the Boltzmann constant, is the volumetric gate dielectric trap density per unit energy, is the oxide capacitance per unit area, is the frequency and is the wentzel-kramers-brillouin (WKB) tunnelling probability for a rectangular barrier. The factor in eq.(3) originates from the integral , being the trap energy and the near-equilibrium occupation function of the defect, which is described by a Fermi function with Fermi level [10]. The integral above is then the area of the bell-shaped function centred at (see Fig.˜5(a)). In other words, only defects with energies within a few of the Fermi level can contribute to the noise. This implies that, as shown in Fig.˜5(a), at K the active energy area is almost two orders of magnitude smaller than at K. On the other hand, the total number of defects showing RTN and, therefore, contributing to LFN, increases dramatically at cryogenic temperatures (see Fig.˜5(b)). One possible way to reconcile these results is to postulate that the defect density increases at low temperatures. As already stressed, at V, most of active defects are in the layer. Hence, the following discussion focuses on the physical origin of electron trapping in this material. Several works [44, 45, 46, 47] showed that electron trapping in is related to disorder, which creates electron traps associated with low-coordinated sites and elongated bonds. The trapping is followed by significant polaron-like distortion of the surrounding amorphous network. At cryogenic temperatures, due to the limited thermal energy available, an electron has a significantly higher probability of remaining near the precursor site in the near-interface oxide long enough to polarize the oxide network, resulting in a shallow trapping event. Afterwards, the electron can either tunnel into this self-created trap state or remain nearby without contributing to the channel current.
@@
As already discussed, at V, is primarily associated with defects due to shift in the active energy area swept by the Fermi level. The physical origin of the kinetics of these defects, which do not freeze out, is still under debate. Recently, in was shown that a full quantum-mechanical solution of the non-radiative multi-phonon (NMP) model can interpret the experimental results at K [19, 9, 48]. According to this solution, nuclear tunnelling becomes the predominant mechanism of charge trapping at cryogenic temperatures. Beside that, the number of detected defects in the layer at higher overdrives is weakly temperature dependent (see Supplement Material), which is challenging to reconcile with the previous discussion based on the trend shown if Fig.˜5(a). Further analyses are required to comprehend the nature of charge trapping in at cryogenic temperatures.
@@ 2.4 Time constants and step heights correlation
Since determines the defect amplitude, if charge trapping processes are dominated by elastic tunnelling, there must be a correlation between and . Many works already proved the inadequacy of elastic tunnelling and, therefore, of McWhorter model to describe LFN in MOSFET devices at K [20, 50, 51]. Here, the possible correlation between and is investigated from K down to K. It is worth to stress that, since the depends with the trap position and, according to eq.(4), depends with the trap position, a linear correlation on a semi-logarithmic axis is expected. In 6(a) and 6(b) the scatter plots at K are reported, together with the calculated Pearson coefficient . No correlation is found, in agreement with [20, 50, 51]. Decreasing temperature, results are similar (see Supplement Materials), down to K (6(c) and 6(d)). From these results it is evident that McWhorter model is not accurate in describing LFN and that defect kinetic is not dominated by the elastic tunnelling, neither at cryogenic temperatures.
@@ 3 Conclusion
The first part of the paper was devoted to the analysis of the threshold voltage shifts due to charged defects, reported in cumulative plots. In this context, it was shown that, at K, the CCDF followed a monomodal exponential distribution, and only defects associated with layer contributed to RTN. At cryogenic temperatures, instead, the CCDF changed its modality with the appearing of other two modes: a small one, attributed to defects, and a large one whose origin is still controversial. An interpretation of both the trimodal shape of the CCDF and the nature of the third branch was provided exploiting the percolation theory. The change in the shape of the CCDF at cryogenic temperatures was attributed to a major sensitivity of the inversion carriers to channel inhomogeneities. Indeed, even though random dopants and surface potential fluctuations are the main cause of the exponential distribution of the already at K, at low temperature carriers’ thermal energy is reduced and they are more susceptible to channel disorder. This larger non-uniformity can induce the formation of insulating regions in the channel, increasing the electrostatic impact of oxide and interface defects: this could origin the third mode of large experimentally observed. Nonetheless, the primary outcome of the first part of this study is the identification of electron trapping within the layer as the dominant contributor for over 80% of the total RTN measured at K.
The second half of this work is dedicated to low frequency noise. The aim is to investigate the contribution of individual defect branches to the total 1/f noise, reconstructed using the original time traces. Even though most of active defects at cryogenic temperatures are associated with the layer, their electrostatic impact is quite small if compared to interface traps. Therefore, most of low frequency noise arises from interfaces defects. Few defects cause most of the 1/f noise at cryogenic temperatures. Nevertheless, LFN generated by bulk traps at K is still comparable to the total LFN at K. An interpretation of this huge activity of bulk defects was proposed and it was based on the concept of electron self-localization due to the polarization of the network. With increasing overdrive voltage, the noise contribution from traps steadily diminishes until it becomes negligible.
@@ 4.1 Smart Array
The chip design also incorporates sense lines at the output of the shift register, as well as on the gate and drain of a device within the array. This allows us to verify the functionality of the periphery and identify potential voltage drops on the gate and drains. Moreover, all measurements are performed using a custom-built switching matrix with separately controllable voltage sources to manage the periphery, along with K2600 SMUs for currentăreadouts.
4.2 Extended Log Likelihood formulation for unknown amount of left censored steps
The regularized formulation of the Likelihood function involves a Multinomial prefactor in terms of the number of steps hypothesized. In many maximum likelihood fitting contexts, the sample size is a constant in terms of the distribution parameters fitted, so that it can be dropped from the maximization. When an unknown amount of left censored steps needs to be included in the fit, however, this prefactor is not a constant anymore. The more generic expression for the Log Likelihood function in terms of the (multimodal) step distribution with PDF needs to be considered then:
with PDF
The binomial prefactor in terms of the unknown total amount of steps has to be absorbed in the Log Likelihood expression to be maximized:
4.3 Suppress over-fitting through constrained maximum likelihood fitting over 6 temperatures simultaneously
By performing a constrained ML fit over the 6 temperatures simultaneously, the impact of over-fitting can be suppressed: imposing a monotonously descending – which is an extra assumption of course – dependence for each mode where it is observable drives the fit towards more realistic values for the distribution parameters.
4.4 Identifying presence of extra step size distribution modes using outlier detection hypothesis test
Starting from monomodal Exponential distribution fits, the presence of extra observable step size distribution modes for each of the CCDF curves fitted can be assessed using an outlier testing hypothesis test, either for the independent unconstrained ML fits, either for the constrained ML fits onto the combined CCDF curves at each of the 6 temperatures measured. The probability of finding the larger value in the sample observed given the distribution modality and distribution parameters estimated can be applied as a significance test:
To avoid subtractive cancellation, it’s safer to compute the tail of the distribution fitted in terms of the test significance probability, and to use the matrix exponential function to maintain precision:
As soon as the tail probability exceeds the critical value for the significance probability chosen, should be considered an outlier for the distribution hitherto fitted. When adhering to the hypothesis of underlying exponential distributions, the presence of an extra exponential mode can then be inferred as an alternate hypothesis, whereupon the ML fit can be repeated accordingly. This process can then be repeated after a bimodal fit to detect the presence of a third (i.c. typically larger) step mode.